OpenStudy (anonymous):

The average annual salary of the employees of a company in the year 2005 was $80,000. It increased by the same factor each year and in 2006, the average annual salary was $88,000. Let f(x) represent the average annual salary, in thousand dollars, after x years since 2005. Which of the following best represents the relationship between x and f(x)? f(x) = 88(0.88)x f(x) = 88(1.1)x f(x) = 80(0.88)x f(x) = 80(1.1)x

OpenStudy (kropot72):

The equation to model the growth is of the form: \[\large f(x)=80000(1+r)^{x}\] where r is the growth factor. We are given that when x = 1, f(x) = 88000 therefore we can write the following equation: \[\large 88000=80000(1+r)\ .........(1)\] Now you need to solve equation (1) to find the value of r.

OpenStudy (anonymous):

88000=80000+80000r =8000

OpenStudy (kropot72):

88000 = 80000 + 80000r subtract 80000 from both sides to get: 8000 = 80000r .........(2) Now divide both sides of (2) by 80000 to get the value of r.

OpenStudy (kropot72):

r = 8000/80000 = ?

OpenStudy (anonymous):

10

OpenStudy (kropot72):

No. Your result is for 80000/8000. You need to calculate 8000/80000 = ?

OpenStudy (anonymous):

0.1

OpenStudy (kropot72):

Correct. r = 0.1. So if we go back to the very first equation to model the growth and plug in the value of r, we get; \[\large f(x)=80000(1+0.1)^{x}\] So if we express the average salary in 2005 in thousands of dollars it becomes 80 and we get: \[\large f(x)=80(1.1)^{x}\]

OpenStudy (anonymous):

thanks

OpenStudy (kropot72):

You're welcome :)